# POPoS: Improving Efficient and Robust Facial Landmark Detection with Parallel Optimal Position Search

### v1

@@ Introduction
Evaluation Metric. Following standard practice in facial landmark detection (Bao et al. 2023), the Normalized Mean Error (NME) is used as evaluation metric, which is defined as $\frac{1}{N_{k}}\sum_{i=1}^{N}\frac{\|\beta_{i}-\hat{\beta}_{i}\|_{2}}{d}$, where $N_{k}$ is the number of landmarks, $\beta$ and $\hat{\beta}$ represent the ground truth and predicted landmarks, respectively, and $d$ is the normalization factor, typically the interpupillary distance.
Implementation Details
Our proposed POPoS framework builds upon the KeyPosS and MMpose [^1] https://github.com/open-mmlab/mmpose frameworks. Under same condition with KeyPosS’s preprocessing and augmentation techniques, we evaluate the performance at varying heatmap resolutions: $64\times 64$, $32\times 32$, $16\times 16$, $8\times 8$, and $4\times 4$ pixels. The backbone network is a lightweight variant of HRNet (Sun et al. 2019), which is pre-trained on ImageNet (Russakovsky et al. 2015) and fine-tuned using the Adam optimizer (Kingma and Ba 2015) with a linear-step decay learning rate schedule. The initial learning rate is set at 2.0e-3, and gradually reducing to 1.0e-5 over 100 epochs. The training is performed on a 4 × NVIDIA 3090 GPU server, with a batch size aligned with KeyPosS (Bao et al. 2023).
Figure 3: Keypoint detection accuracy at heatmap resolutions of $64\times 64$, $32\times 32$, $16\times 16$, $8\times 8$, and $4\times 4$. (The smaller radar map denotes the better performance)
@@
STAR Loss (Zhou et al. 2023b) 4.02 2.87 - 4.62 13.37M -
@@
Comparison with State-of-the-Art Methods
@@
Comparison with Current Decoding Schemes
The core problem addressed in this article is the verification of decoding accuracy at different resolutions. The decoding effects of five different methods were compared across five datasets under various heatmap resolution settings. As shown in Fig. 3, a distinctive advantage of POPoS is found in the integration of the Iterative Gauss–Newton Optimization (IGNO) approximation method with a strategic sampling approach. This integration is especially beneficial in low-resolution settings, where high-quality results are delivered efficiently by POPoS.
@@
Figure 5: Detailed analysis of POPoS performance factors: (a) Effect of sampling area (b) Effect of sampling rate $\tau$ (c) Anchor number and (d) Local distance loss weight, each illustrating their impact on the accuracy of predictions
Figure 6: Visualization of POPoS’s performance in low-resolution settings: Comparative analysis with one-hot, two-hot, dark-pose, and KeyPosS methods. Ground truth is represented by green dots, predictions by red dots, discrepancies by blue lines, and significant inaccuracies are circled in yellow. (Best viewed in color and at a higher zoom level.)
This part offer an in-depth analysis on the effect of the proposed modules and the hyperparameter settings.
@@
The efficacy of the Iterative Gauss–Newton Optimization (IGNO) for pseudo-range multilateration (PRM) was examined by integrating it into the baseline model, replacing the least squares approximation. Iterations were set to a maximum of 20 or terminated upon convergence. The KeyposS (Bao et al. 2023) was considered as the baseline. Notable improvements were observed in the 300W and COFW datasets as shown in Fig. 4. This outcome is attributed to a decrease in prediction errors with a larger number of anchors.
@@
The influence of Multilateration Anchor (MA) loss in addressing encoding errors was evaluated by retraining the HRNet network across various resolutions. The results showed significant improvements of the POPoS in all datasets when the multilateration anchor loss was integrated, as shown in Fig. 4. Specifically, in the 4x4 resolution heatmap, the prediction errors decreased by 47% for the 300W dataset and 58% for the COFW dataset.
@@
The Potential Position Parallel Sampling and Computing (PPPSC) is tested on five datasets at different resolutions, as shown in Fig. 4. This strategy notably bolstered the stability and performance of the IGN Optimization algorithm. For the ablation experiments on the parameter settings of sampling range $(h,w)$ and sampling frequency $\tau$, tests are conducted on the 300W dataset to evaluate the impact of different parameters on the experimental results at a resolution of $32\times 32$ resolution heatmap. The experimental outcomes are illustrated in Fig. 5. Ablation studies on the sample area indicated that the optimal values are distributed within the range of 1-2 pixels around the maximum value. Increasing the sampling range does not improve decoding accuracy. This suggests that, for certain challenging samples, errors already exist during the model’s prediction phase, specifically in the encoding stage, and cannot be optimized by reducing decoding errors. Ablation studies on the sample rate $\tau$ indicated that optimal performance is attained with approximately 10 sampling points per pixel length, equating to about 1.2 points per original image pixel as shown in Fig. 5a. Further increasing the number of sampling points yielded negligible improvements.
@@
The ablation study of anchor number $N_{a}$ in the decoding step, as illustrated in the Fig. 5c indicating that too few anchor results in inaccurate predictions, while excessive anchor introduces larger calculation error. In the case of a $32\times 32$ resolution heatmap, the optimal number of anchors is determined to be 25. An ablation experiment on the weight of the local anchor loss in the 300W dataset (Fig. 5d) demonstrated consistent improvement in prediction accuracy across various weight ratios, with the most effective being a ratio of 6. This finding underscores the critical impact of encoding accuracy errors on pseudo-range multilateration and the effectiveness of the local anchor loss in mitigating convergence issues.
Visualization Analysis
To evaluate POPoS’s adaptability in practical applications, a comprehensive visualization analysis was conducted. The results were compared against traditional methods such as one-hot, two-hot, DarkPose, and KeyPosS. This comparison focused on challenging low-resolution scenarios, as depicted in Fig. 6. The analysis highlighted that traditional methods like one-hot, two-hot, and dark-pose struggle with precision at lower heatmap resolutions such as $4\times 4$ and $8\times 8$. POPoS shows superior performance at $4\times 4$ and higher resolutions. Furthermore, POPoS consistently addresses localization errors across all examined resolutions. This proficiency is particularly evident in Fig. 6, where POPoS significantly outperforms the traditional methods.
The visualization confirms the robustness of POPoS, demonstrating its capability to optimize encoding and decoding errors effectively in diverse application scenarios.
Conclusion
In summary, a noval heatmap based method with accuracy-efficiency trade-offs named Parallel Optimal Position Search (POPoS) framework is built. In encoding stage, the multilateration anchor loss function is proposed, which reduce the prediction error of heatmap. In the decoding stage, an optimization-based multilateration method is employed, which overcomes accuracy constraints in traditional method. And the Equivalent Single-Step Parallel Computation algorithm is proposed to significantly enhance computational efficiency. Our testing results on five widely recognized datasets demonstrate that POPoS performs exceptionally well, particularly under low-resolution heatmap conditions. In general, POPoS achieves effect balance between accuracy and efficiency. This advancement demonstrates a significant leap forward in the facial landmark detection domain.

### v5

@@ 4.1 Dataset and Metric
4.2 Evaluation Metric
Following standard practice in facial landmark detection (Bao et al. 2023), the Normalized Mean Error (NME) is used as evaluation metric, which is defined as $\frac{1}{N_{k}}\sum_{i=1}^{N}\frac{\|\beta_{i}-\hat{\beta}_{i}\|_{2}}{d}$, where $N_{k}$ is the number of landmarks, $\beta$ and $\hat{\beta}$ represent the ground truth and predicted landmarks, respectively, and $d$ is the normalization factor, typically the interpupillary distance.
4.3 Implementation Details
Our proposed POPoS framework builds upon the KeyPosS and MMpose [^1] https://github.com/open-mmlab/mmpose frameworks. Under same condition with KeyPosS’s preprocessing and augmentation techniques, we evaluate the performance at varying heatmap resolutions: $64\times 64$, $32\times 32$, $16\times 16$, $8\times 8$, and $4\times 4$ pixels. The backbone network is a lightweight variant of HRNet (Sun et al. 2019), which is pre-trained on ImageNet (Russakovsky et al. 2015) and fine-tuned using the Adam optimizer (Kingma and Ba 2015) with a linear-step decay learning rate schedule. The initial learning rate is set at 2.0e-3, and gradually reducing to 1.0e-5 over 100 epochs. The training is performed on a 4 × NVIDIA 3090 GPU server, with a batch size 32. In the decoding stage of POPoS, the key hyperparameters, such as the sampling range (w=1) and the number of anchors (K=10), are determined through training on the 300W (Sagonas et al. 2016) dataset.
Figure 3: Keypoint detection accuracy at heatmap resolutions of $64\times 64$, $32\times 32$, $16\times 16$, $8\times 8$, and $4\times 4$. [The smaller radar map denotes the better performance]
@@ 4.3 Implementation Details
STAR Loss (Zhou et al. 2023) 4.02 2.87 - 4.62 13.37M -
@@
4.4 Comparison with State-of-the-Art Methods
@@
4.5 Comparison with Decoding Schemes
The key problem addressed in this article is the verification of decoding accuracy at different resolutions. The decoding effects of five different methods were compared across five datasets under various heatmap resolution settings. As shown in Fig. 3, a distinctive advantage of POPoS is found in the integration of the Iterative Gauss–Newton Optimization (IGNO) approximation method with a strategic sampling approach. This integration is especially beneficial in low-resolution heatmaps settings, where high-quality results are delivered efficiently by POPoS.
@@ 4.5 Comparison with Decoding Schemes
Figure 5: Detailed analysis of POPoS performance factors: (a) Effect of sampling area (b) Effect of sampling rate $\tau$ (c) Anchor number and (d) MA loss weight, each illustrating their impact on the accuracy of predictions.
Figure 6: Visualization of POPoS’s performance in low-resolution heatmaps settings: Comparative analysis with one-hot, two-hot, dark-pose, and KeyPosS methods. Ground truth is represented by green dots, predictions by red dots, discrepancies by blue lines, and significant inaccuracies are circled in yellow. [Best viewed in color and at a higher zoom level]
4.6 Ablation Studies
This section offers an in-depth analysis of the effects of the proposed modules and hyperparameter settings.
@@ 4.6 Ablation Studies
The effectiveness of Iterative Gauss–Newton Optimization (IGNO) for pseudo-range multilateration (PRM) was evaluated by integrating it into the baseline model, replacing the least squares approximation. The number of iterations was capped at 20 or halted upon convergence. KeyPosS (Bao et al. 2023) served as the baseline. Improvements were observed in the 300W and COFW datasets, as illustrated in Fig. 4. These enhancements are attributed to a reduction in prediction errors achieved through an increased number of anchors.
@@
The impact of Multilateration Anchor (MA) loss on mitigating encoding errors was assessed by retraining the HRNet network across various resolutions. The integration of the multilateration anchor loss resulted in significant improvements in POPoS across all datasets, as depicted in Fig. 4. Specifically, in the $4\times 4$ resolution heatmap, prediction errors decreased by 47% for the 300W dataset and by 58% for the COFW dataset.
@@
Potential Position Parallel Sampling and Computing (PPPSC) was evaluated on five datasets at varying resolutions, as presented in Fig. 4. This approach significantly enhanced the stability and performance of the IGN Optimization algorithm. Ablation experiments on the sampling range $(h,w)$ and sampling frequency $\tau$ were conducted using the 300W dataset at a resolution of $32\times 32$. The results, shown in Fig. 5, indicated that the optimal sample area values lie within 1-2 pixels around the maximum value. Increasing the sampling range did not improve decoding accuracy, suggesting that some errors originate in the encoding stage and cannot be corrected by decoding adjustments. Additionally, ablation studies on the sample rate $\tau$ demonstrated that optimal performance is achieved with approximately 10 sampling points per pixel length, equivalent to about 1.2 points per original image pixel, as shown in Fig. 5a. Further increases in the number of sampling points resulted in negligible improvements.
@@
The ablation study on the number of anchors $N_{a}$ in the decoding step, as shown in Fig. 5c, indicates that too few anchors result in inaccurate predictions, while an excessive number of anchors introduces larger calculation errors. For a $32\times 32$ resolution heatmap, the optimal number of anchors was determined to be 25. Additionally, an ablation experiment on the weight of the local anchor loss in the 300W dataset (Fig. 5d) demonstrated consistent improvements in prediction accuracy across various weight ratios, with the most effective ratio being 6. This finding highlights the significant impact of encoding accuracy errors on pseudo-range multilateration and the efficacy of the local anchor loss in mitigating convergence issues.
4.7 Visualization Analysis
To evaluates POPoS’s adaptability in practical applications, we conducted a comprehensive visualization analysis comparing it with traditional methods such as one-hot, two-hot, DarkPose, and KeyPosS. Focusing on challenging low-resolution heatmaps scenarios depicted in Fig. 6, the analysis revealed that conventional methods struggle with precision at heatmap resolutions of $4\times 4$ and $8\times 8$. In contrast, POPoS demonstrates superior performance starting from $4\times 4$ and maintains high accuracy at higher resolutions. Additionally, POPoS effectively mitigates localization errors across all tested resolutions, as clearly shown in Fig. 6. These visualizations confirm POPoS’s robustness, highlighting its ability to optimize encoding and decoding processes effectively across diverse application scenarios.
5 Conclusion
We introduce Parallel Optimal Position Search (POPoS), a heatmap-based framework that effectively balances accuracy and efficiency in facial landmark detection (FLD). POPoS employs the Multilateration Anchor Loss during encoding to reduce heatmap prediction errors by enhancing anchor point accuracy. In decoding, an optimization-based multilateration method surpasses traditional approaches in landmark localization. Additionally, the Equivalent Single-Step Parallel Computation algorithm significantly improves computational efficiency, enabling real-time FLD applications. Extensive evaluations on five benchmark datasets demonstrate that POPoS consistently outperforms existing FLD methods, particularly in low-resolution scenarios.
Acknowledgments
Zhi-Qi Cheng’s research for this project was supported in part by the US Department of Transportation, Office of the Assistant Secretary for Research and Technology, under the University Transportation Center Program (Federal Grant Number 69A3551747111). This work was also supported in part by the National Natural Science Foundation of China (Grant No. 62372387) and the Key R&D Program of Guangxi Zhuang Autonomous Region, China (Grant Nos. AB22080038 and AB22080039).